H* weekly meeting
- THz sensors: Fredrik ask for a quote Gentec +SLT
- Fredrik talks to J-F for the filters
- Lamp arrived at B275 : look for the doc and get cables, check how igniter works ->Maxime
- Tim simulations: was a bug with polarization - recovered what we had previously. Effect of smaller BW is small. 5% BW still gets 60% in GS.
Look into having the code running with n=30 maybe on a VM?
- Reminder: Think about the question of high rep, femtosecond laser
- H* BEAMLINE
- KF25 blind flange -> drill
- Teflon o-ring: Chloé will ask the cryolab, Maxime follows up with Martin , KF16 blind
- Fredrik found a CF40 gate valve. Find a solution
- adapt the telescope outside the laser to adjust for the height
- need to double check if pressure's fine with the electrode
- install Helmholz coils and measure the field with 1 fluxgate
- install MCP feedthroughs